OpenStudy (inowalst):

My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years ?

OpenStudy (calculusxy):

g= grandson's age s= son's age m = grandfather's age "My grandson is as many days as my son in weeks" s=7g "My grandson is as many months as I am in years" m+12g We know that: g + s + m = 120 So we can substitute s and m: g + 7g + 12g = 120 20g = 120 g= 6yrs as of the grandson's age Now do: s = 7(6) = 42 yrs as of son's age m = 12(6) = 72yrs as of my age (grandpa's age)

OpenStudy (anonymous):

You are 120*12 / (1 + 7 + 12) yrs old = 72 yrs old grand son 72 months = 6 years son 7*6yrs = 42 yrs old. 7 days in the week, and 12 months in the year Sound about right?

OpenStudy (inowalst):

@calculusxy Thank you! @BasicallyAwesome Thank you, as well.

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (anonymous):

Anytime :)

OpenStudy (theraggedydoctor):

dot.